# Memory-Centric Embodied Question Answer

### v1

@@ 3.2.1 Memory Storage
The memory information is stored in dense vectors to facilitate efficient retrieval. We employ a unified encoder $e_{c}(\cdot)$ to extract multi-modal memory features $f_{i}=e_{c}(M_{i})$. These feature vectors are then stored in a vector library =$\mathcal{L}=\{L_{i}\}_{k=0}^{N_{s}}=\{<I_{i},M_{i},f_{i}>\}_{i=0}^{N_{s}}$, with each memory assigned a unique index $I_{i}$ for subsequent retrieval operations.
3.2.2 Memory Retriever
Before the reasoning process within each module, it is essential to retrieve the necessary memories from the memory database. To this end, we have designed a specialized retriever tailored for the hierarchical memory retrieval system. Our memory retriever $\mathcal{F}_{r}$ consists of two components:
Scene Retrieval.
Scene retrieval occurs only when the agent initializes. Upon obtaining the initial observation $o_{0}$, we extract its feature $f_{o_{0}}=e_{c}(o_{0})$. We compute cosine similarity between $f_{o_{0}}$ and vectors $\mathcal{L}$ in the vector library, and select the $\text{top}_{k}$ most similar features (requiring similarity scores to exceed a threshold)
$I^{s}=\text{Index}_{\mathcal{L}}(\text{top}_{k}(\text{sim}(f_{o_{0}},\mathcal{L}))),$ (2)
where $I^{s}$ is the set of scene index, and $\text{Index}_{\mathcal{L}}(\cdot)$ is the operator of index in $\mathcal{M}$. Next, we identify the scene index $I^{c}=\text{Max}(\text{Fre}(I^{s}))$ with the highest number of occurrences as the current scene, $\text{Fre}(\cdot)$ is the operator used to calculate the frequency of occurrence of elements. The vectors associated with this scene form a sub-database for subsequent retrieval steps.
Content Retrieval.
@@ 3.2.2 Memory Retriever
For content retrieval, we adopt a multi-modal retrieval approach. Given the current observation $o_{i}$ and user question $Q$, we use a unified encoder to generate a combined feature
$f_{q}=\text{Norm}(\operatorname{concat}(e_{c}(o_{i}),e_{c}(q))).$ (3)
We then compute similarity between $f_{q}$ and vectors in the scene memory $M_{i}$ (from the result of scene retrieval and prior exploration) and retrieve the $\text{top}_{k}$ most similar indices as the required memory content. We final get the local memory
$\begin{split}\mathcal{M}^{s}=\text{top}_{k}(\{M_{i}|\|(f_{q}-M_{i}\|_{2}<\alpha_{e}(1+\text{Entropy}(f_{q})))\cap\text{sim}(f_{q},M_{i})>\alpha_{s}\}),\end{split}$ (4)
where $\mathcal{M}^{s}$ represents the memory required in the current state, $\alpha_{e}$ and $\alpha_{s}$ are the weight of similarity. We utilize entropy scaling to adapts the search scope to query complexity and semantic similarity greater than a threshold $\alpha_{s}$.
@@
The size of memory significantly impacts model performance (see supplementary material). To address this, we employ a dynamic k-value mechanism that adaptively samples varying amounts of memory information according to different scenarios. Specifically, we implement an entropy-based linear adjustment strategy formulated as:
@@
[…]
3.2.3 Memory Injector
The injection of global memory only needs to be loaded before the agent makes its first step of exploration. For local memory, for the planner, we map $n$ candidate exploration directions to pixel space, marking the candidate directions in the current observed RGB image to obtain $d_{i}$. We use $d_{i}$, memory, and the question as inputs and prompt the VLM to select one of the candidate directions. Based on the VLM’s choice, we construct a one-hot vector as the semantic weight for the planner. For the answering module, we directly input the current observation $o_{i}$, memory, and the question as prompts to the VLM to obtain the response.
3.2.4 Memory Update
After loading the global memory, continued exploration may reveal new environmental changes that necessitate memory updates. To address this, we have designed a set of memory update mechanism to effectively prevent erroneous triggering of updates. The specific rules are as follows:
  • Agent enters novel region with
    $\min_{m_{j}\in M}\|P_{i}-P_{j}\|_{2}>\beta_{p},$ (6)
    $\min_{m_{j}\in M}\angle(R_{i},R_{j})>\beta_{r},$ (7)
    where $M$ is the set of memories. $P_{i}$ and $R_{i}$ is the $i$-th position and rotation in memories. $\|\cdot\|_{2}$ is the Euclidean distance operator. (,$\angle$($\cdot$,$\cdot$)∠ ( ⋅ , ⋅ ) is the Euclidean angle operator. In other words, the distance between the current position and orientation and all the positions in memory is greater than threshold $\beta_{p}$ and $\beta_{r}$.
  • The similarity between the current observation and the memory observation is below a threshold. We calculate similarity
    by integrating both structural similarity and semantic similarity, where $(f_{i},f_{j})=(e_{c}(o_{i}),e_{c}(o_{j}))$ is current observation feature and history observation feature respectively. The SSIM [33] is structural similarity.
The memory update mechanism is triggered only when all the above rules are satisfied. Due to the nature of our memory storage, updating memory can be efficiently achieved by modifying the corresponding memory index.
3.3 Discrete Stop Criterion
We adopt a discrete stopping criterion, where we set a multiple-choice question and let the model assess its confidence level in answering the question. If the confidence exceeds a given threshold, the exploration stops, and the answering module is invoked. If the maximum number of steps is reached, the agent will also stop exploring and provide an answer. So the probability modeling of the stop criterion is $\mathcal{F}_{s}=\mathbf{1}(\mathcal{P}(S|M_{i})>\gamma),$ where $\mathbf{1}$ is an indicator function, which is 1 when the condition is met and 0 otherwise. $\gamma$ is threshold, and $S$ means current stats of the agent. In the implementation, we prompt VLMs to provide the confidence level in answering the question based on the retrieved memory information from the current observation. For detailed information, please refer to the supplementary material.
@@ 3.4 Answering Module
We employ a retrieval-augmented VLM as the answering module, which is invoked after the agent terminates exploration. By selectively feeding only memory content relevant to the query into the VLM, we achieve both computational efficiency and reduced interference from extraneous information.
4 Data Construction
To evaluate memory ability of the models, we propose an multi-target EQA dataset based on HM3D [26]. Our dataset includes 1587 QA samples on 500 scenes (more statistical data see the supplemental material). The most significant distinction from existing datasets lies in the fact that the MT-HM3D dataset constructs question-answering pairs targeting multiple regions and multiple objects in real-world scenarios, while exhibiting a richer diversity of question formats. We define four types of questions:
  • Comparison: comparing status two or more entities, e.g., “Is the bed closer to the chair than to the window in the bedroom?”
@@ 4 Data Construction
We use GPT-4o [23] and GPT-4-mini [23] to construct multi-target question answering. The process consists of the following steps: (1) Sample 5 to 10 images from the scene using the Habitat simulator. (2) Input the sampled images and prompt the GPT-4o [23] select multiple related objects from those images, specifying the output format to obtain raw data. (3) Use GPT-4-mini [23] to filter the raw data, removing illogical question-answer pairs, and generate diverse QA data. (4) Manually verify the correctness of the data to ensure that the questions can be answered within the embodied environment. Specific prompts and more details are provided in the supplementary material.
Metrics. We evaluate the embodied question-answering model using success rate (%), normalized steps, GPT score, and ROUGEL [19] (the details see supplemental material).
@@ 5.1 Implementation
We use QwenVL2-7B [32] and GPT-4o [23] as the built-in multi-modal large models, and employ Clip-ViT-Large [25] as unified encoder to encode the multi-modal memory into 768-dimensional features. The object detector is YOLOv11[18]. Our simulator follows the setup of Explore-EQA [27] and conducts experiments in Habitat-Sim [30], and we conduct real-world experiments using the Unitree Go2 robot (see supplementary material). Inference is performed on an NVIDIA Tesla L40, with approximately 40 hours required to complete inference on the MT-HM3D dataset. We use Faiss [13] as a library of dense vectors.
Table 1: Experiments on multiple methods across MT-HM3D, HM-EQA, and OpenEQA, testing various foundational models. denotes that the value was converted based on the metrics provided in the original paper. $\ddagger$ indicates that the performance was obtained through our replication.
@@
1 ExploreEQA Prismatic-7B 35.1 0.64 51.5 0.75 28.28 0.26 0.695
2 Qwen2VL-7B 36.1 0.66 44.31 0.74 27.36 0.241 0.741
3 GPT4o 36.2 0.64 47.4 0.77 29.21 0.255 0.773
4 GraphEQA Qwen2VL-7B 45.63 0.45 56.7 0.35 30.1 0.277 0.625
5 GPT4o - - 63.5 0.2 - - -
6 MemoryEQA Qwen2VL-7B 51.2 0.4 59.9 0.52 34.64 0.3 0.6
7 GPT4o 55.1 0.41 63.4 0.4 36.8 0.32 0.552
@@ 5.2 Main Results
As shown in Table 1, we evaluate three approaches ExploreEQA [27], GraphEQA [29], and MemoryEQA across the MT-HM3D, HM-EQA [27], and OpenEQA [21] datasets, testing their compatibility with diverse VLMs (Prismatic [17], Qwen2VL [32] and GPT4o [23]).
On MT-HM3D, MemoryEQA attains a success rate of 54.5%, outperforming baseline by 18.9% (Exp.3), highlighting the critical role of hierarchical memory in multi-target tasks. The results demonstrate that MemoryEQA exhibits superior performance in multi-modal reasoning tasks, particularly in complex scene understanding and knowledge integration.
The analysis of experimental results across Exp.1-3, Exp.4-5, and Exp.6-7 reveals a significant positive correlation between the performance of the VLM and the effectiveness of the EQA system. This observation underscores the critical role that VLM plays in enhancing the EQA system’s ability to process and interpret complex queries within visual environments. As the VLM’s accuracy and understanding improve, so does the EQA system’s capacity to deliver precise and contextually relevant answers, demonstrating a synergistic relationship between the two components. This finding highlights the importance of advancing VLM capabilities to further boost the overall performance of EQA systems in practical applications.
Table 2: Ablation study on MT-HM3D and HM-EQA to evaluate the impact of injecting memory into different modules on performance. (S: Stop Criterion, A: Answering Module, P: Planner)
@@
None 34.3 0.7 44.6 0.79
S 35.1 0.66 43.5 0.7
S + A 46.8 0.69 52.2 0.71
S + A + P 51.2 0.4 59.9 0.52
Table 3: The impact of global memory and local memory on MemoryEQA.
Method Succ. $\uparrow$ Step $\downarrow$
ExploreEQA 51.5 0.75
MemoryEQA w/o Global Memory 56.4 0.74
MemoryEQA w/o Local Memory 52.15 0.6
MemoryEQA 59.9 0.52

### v2

@@ 3.2.2 Memory Module
Storage. The memory information is stored in dense vectors to facilitate efficient retrieval. We employ a unified encoder $e_{c}(\cdot)$ to extract multi-modal memory features $f_{i}=e_{c}(m_{i})$. These feature vectors are then stored in a vector library $\mathcal{L}=\{L_{i}\}_{k=0}^{N_{t}}=\{<I_{i},p_{i},r_{i},m_{i},f_{i}>\}_{i=0}^{N_{t}}$ with each memory assigned a unique index $I_{i}$ for subsequent retrieval operations. In addition, once the agent completes a task, the memory accumulated in that scene is persistently stored in scene memory library $\mathcal{M}^{s}=\{\mathcal{M}_{t}\}^{N_{s}}_{t=0}$, where $N_{s}$ is the number of scenes. It serves as the initial memory repository for future tasks performed in the same environment.
Update. During the exploration process, the agent may reveal new environmental changes that necessitate memory updates. So we design a memory update mechanism to effectively prevent erroneous triggering of updates. The specific rules are as follows:
  • We check whether the agent enters a known region with
    $\min_{m_{j}\in\mathcal{M}}\|p_{i}-p_{j}\|_{2}<\beta_{p},$ (3)
    $\min_{m_{j}\in\mathcal{M}}\angle(r_{i},r_{j})<\beta_{r},$ (4)
    where $\|\cdot\|_{2}$ is the Euclidean distance operator. $\angle(\cdot,\cdot)$ is the Euclidean angle operator. In other words, if the distance between the current position and all positions in memory is greater than the threshold $\beta_{p}$, and the angle between the current orientation and all orientations in memory is greater than the threshold $\beta_{r}$, we regard that the agent enters a new region.
  • The similarities between the current observation and all memory observations exceed a threshold. We calculate the similarity
    by integrating both structural similarity and semantic similarity, where $(f_{i},f_{j})=(e_{c}(o_{i}),e_{c}(o_{j}))$ denotes the current observation feature $f_{i}$ and the history observation feature $f_{j}$, respectively. The SSIM [wang2004image] is structural similarity, and measures the perceptual similarity between two images by comparing their luminance, contrast, and structural patterns. The $\text{sim}(\cdot)$ is the cosine similarity.
The memory update mechanism is triggered if and only if the three rules above are satisfied. Otherwise, the observation is regarded as entering a new region, and the memory is directly added to the memory library. Once an update is triggered, we assign the index $I_{i}$ of the memory to be updated to the new memory, and then remove the original memory to complete the update process.
@@
Retrieval. We adopt a multi-modal retrieval approach. Given the initial observation $o_{0}$ at a scene and the user question $q$, we use the unified encoder $e_{c}(\cdot)$ to generate a combined feature
$f_{q}=\text{Norm}(\operatorname{concat}(e_{c}(o_{0}),e_{c}(q))).$ (6)
Then, we use $f_{q}$ as the query to retrieve memory information related to the current scene from the scene memory library $\mathcal{M}^{s}$ via cosine similarity, forming the initial memory library $\mathcal{M}$. The above process is performed only once, before the agent begins exploration. At each exploration step, we retrieve from the memory library $\mathcal{M}$ for the memory information $\mathcal{M}^{cur}$ required by different modules. We formalize the retrieval process as
$\begin{split}\mathcal{M}^{cur}_{\{p,s,a\}}=\text{top}_{k}(\{m_{i}|(\|e_{c}(q_{\{p,s,a\}})-e_{c}(m_{i})\|_{2}\\ <\alpha_{e}(1+\text{Entropy}(e_{c}(q_{\{p,s,a\}}))))\\ \cap\text{sim}(e_{c}(q_{\{p,s,a\}}),e_{c}(m_{i}))>\alpha_{s}\}),\end{split}$ (7)
where $\alpha_{e}$ and $\alpha_{s}$ are the weight of similarity, and $\mathcal{M}^{cur}_{\{p,s,a\}}$ denote the memory information for the planner, stopping module, and answering module, respectively. We use entropy to measure the uncertainty of the query features, making it an adaptively adjustable distance threshold. When the entropy is low (i.e., the query semantics are clear), we aim to match only very close memories to avoid noise; when the entropy is high (i.e., the query semantics are ambiguous), we allow a wider matching range, otherwise no relevant memories might be retrieved. Additionally, the cosine similarity between the memory and the query should exceed the threshold $\alpha_{s}$.
@@
The size of $\mathcal{M}^{cur}_{\{p,s,a\}}$ significantly impacts model performance (see supplementary material). To address this, we employ a dynamic k-value mechanism that adaptively samples varying amounts of memory information for the $\text{top}_{k}$ selection according to the complexities of the user query $q$ and the initial observation $o_{0}$. Specifically, we implement an entropy-based linear adjustment strategy for the value of $k$, formulated as
@@
where $k_{min}$ denotes the minimum allowable value of $k$, $\left\lceil\cdot\right\rceil$ is the operator of floor, $\beta$ is the weight of entropy. Intuitively, this means that complex queries and scenes will retrieve more memory information.
Table 1: The comparison between our proposed MT-HM3D and existed dataset. We assume that each question involves at least 1 region, even if the question does not explicitly mention any region.
Dataset Simulator Source Real Scenes Objects Region Creation
MT-EQA[yu2019multi] House3D SUNCG 3.20 1.0 Rule-Based
EXPRESS-Bench[jiang2025beyond] Habitat HM3D 3.21 1.09 VLMs
HM-EQA[ren2024explore] Habitat HM3D 2.87 1.0 VLMs
Open-EQA[majumdar2024openeqa] Habitat ScanNet/HM3D 2.53 1.0 Manual
MT-HM3D (Ours) Habitat HM3D 3.44 1.49 VLMs
Figure 2: The category distribution of MT-HM3D.
3.3 Stopping Criterion
We combine the retrieved memory information with the current observation and prompt a VLM to obtain the current stopping confidence. If the confidence exceeds a given threshold, the exploration stops, and the answering module is invoked. If the maximum number of steps is reached, the agent will also stop exploring, and the answering module is invoked. So the stopping module is formulated as $\mathcal{F}_{s}=\mathbf{1}(\mathcal{P}(s|o_{i},f_{M}^{r}(q_{s}))>\gamma),$ where $s$ is the confidence for stopping, $\mathbf{1}$ is an indicator function, which is 1 when the condition is met and 0 otherwise. $\gamma$ is the threshold. In the implementation, we use a VLM as $\mathcal{P}$ and prompt it to provide the confidence in stopping the exploration process based on the retrieved memory information and the current observation. For detailed information, please refer to the supplementary material.
@@ 3.4 Answering Module
We employ a retrieval-augmented VLM as the answering module, which is invoked once the agent terminates exploration. Memory retrieval mechanism identifies and extracts only the memory content that is semantically relevant to the given question. This selective feeding strategy not only reduces computational overhead but also mitigates the risk of distraction from irrelevant or noisy information. Moreover, by grounding the VLM’s reasoning in the retrieved memory rather than raw observations, the answering process becomes more interpretable and contextually coherent, enabling the model to generate responses that faithfully reflect the agent’s past experiences and the spatial-temporal context of the environment.
4 Benchmark Construction
To evaluate the memory ability of EQA agents, we propose a multi-target EQA dataset based on HM3D [ramakrishnan2021habitat], namely MT-HM3D. Our MT-HM3D dataset includes 1587 QA samples on 500 scenes (more statistical data is in the supplemental material).
As shown in Table 5, the most significant distinction from existing datasets lies in the fact that the MT-HM3D constructs question-answering pairs targeting multiple regions and multiple objects in real-world scenarios, while exhibiting a richer diversity of question formats. We define four types of questions.
  • Comparison: comparing the status of two or more entities, e.g., “Is the bed closer to the chair than to the window in the bedroom?”
@@ 4 Benchmark Construction
The statistical details of the MT-HM3D dataset are illustrated in Figure 4.
We use GPT-4o [gpt4o] and GPT-4-mini [gpt4o] to construct multi-target question answering. The process consists of the following steps: (1) Sample 5 to 10 images from the scene using the Habitat simulator. (2) Input the sampled images and prompt GPT-4o [gpt4o] to select multiple related objects from these images, specifying the output format to obtain unfiltered data. (3) Use GPT-4-mini [gpt4o] to filter the raw data, removing illogical question-answer pairs, and generate diverse QA data. (4) Manually verify the correctness of the data to ensure that the questions can be answered within the embodied environment. Specific prompts and more details are provided in the supplementary material.
Metrics. For multiple-choice questions, we make agents directly provide a predicted option, and then calculate the success rate (%). For open-vocabulary questions, we use two metrics to evaluate the performance of agents: GPT score and ROUGEL [lin2004rouge]. Then we use average normalization steps to evaluate the efficiency of the agents, and the specific calculation formula is
$Norm_{step}=\frac{1}{N}\sum_{i=0}^{N}\frac{N_{i}}{\sqrt{S_{i}}*\gamma_{s}},$ (9)
where $N$ means the number of samples, $N_{i}$ represents the explore step number of agents on the sample, $S_{i}$ is room size, and $\gamma_{s}$ is the relation ratio between max steps and room size.
@@ 5.1 Implementation
We use Qwen30VL-8B [Qwen2.5-VL] and GPT-4o [gpt4o] as the built-in VLMs, and employ Clip-ViT-Large [radford2021learning] as the unified encoder to encode the multi-modal memory into 768-dimensional features. The object detector is YOLOv11[khanam2024yolov11]. We use Faiss [douze2024faiss] as a library of dense vectors. Our simulator follows the setup of Explore-EQA [ren2024explore] and conducts experiments in Habitat-Sim [szot2021habitat]. Inference is performed on a single NVIDIA Tesla L40, with approximately 40 hours required to complete inference on the MT-HM3D dataset.
Table 2: Experiments on multiple methods across MT-HM3D, HM-EQA, and OpenEQA, testing various foundational models. $\dagger$ indicates the performance of the official implementation.
@@
Video LLM
1 ExploreEQA[ren2024explore] VideoMind-7B[liu2025videomind] 35.13 0.52 51.3 0.61 31.43 0.28 0.61
2 InternVideo2.5-8B[wang2025internvideo2] 39.58 0.53 53.52 0.59 33.14 0.3 0.59
3 MemoryEQA VideoMind-7B[liu2025videomind] 41.93 0.46 54.44 0.58 31.34 0.29 0.58
4 InternVideo2.5-8B[wang2025internvideo2] 41.15 0.46 53.36 0.55 33.69 0.31 0.58
Spatial LLM
5 ExploreEQA[ren2024explore] Spatialbot-3B[cai2025spatialbot] 31.53 0.53 49.61 0.71 25.97 0.15 0.61
6 SpatialVLM-7B[chen2024spatialvlm] 33.78 0.55 53.98 0.61 30.5 0.25 0.7
7 MemoryEQA Spatialbot-3B[cai2025spatialbot] 32.01 0.42 51.55 0.51 26.97 0.14 0.55
8 SpatialVLM-7B[chen2024spatialvlm] 40.95 0.41 55.55 0.50 29.33 0.26 0.71
Visual LLM
9 ExploreEQA[ren2024explore] Prismatic-7B[karamcheti2024prismatic] 30.22 0.64 51.5 0.75 28.28 0.26 0.7
10 Qwen2VL-7B[wang2024qwen2] 33.18 0.76 44.31 0.74 27.36 0.24 0.74
11 Qwen3VL-7B[Qwen2.5-VL] 33.43 0.7 50.13 0.69 31.29 0.26 0.71
12 GPT4o[gpt4o] 33.21 0.64 47.4 0.77 29.21 0.26 0.77
13 MemoryEQA Qwen2VL-7B[wang2024qwen2] 40.4 0.4 55.9 0.52 27.64 0.28 0.6
14 Qwen3VL-7B[Qwen2.5-VL] 41.95 0.4 56.65 0.49 30.53 0.3 0.61
15 GPT4o[gpt4o] 43.11 0.41 61.4 0.4 30.87 0.32 0.58
@@ 5.2 Main Results
We compare MemoryEQA with ExploreEQA [ren2024explore] using three types of VLMs (Video LLMs [liu2025videomind, wang2025internvideo2], Spatial LLMs [cai2025spatialbot, chen2024spatialvlm], and Multi-Image LLMs [wang2024qwen2, Qwen2.5-VL, gpt4o, karamcheti2024prismatic]) and evaluate them on MT-HM3D, HM-EQA [ren2024explore], and OpenEQA [majumdar2024openeqa].
On MT-HM3D, MemoryEQA attains a success rate of 43.11%, outperforming baseline by 9.9% (Exp.12 vs Exp.15), highlighting the critical role of hierarchical memory in multi-target tasks. By comparing Exp.1-2 and Exp.9-12, it can be observed that the VideoLLM-based EQA method shows an improvement on MT-HM3D (ranging from 1.92% to 9.36%), as the video contains historical information. However, due to the excessive presence of irrelevant information within the video memory, which interferes with the final answer, the performance is suboptimal. Comparing Exp.1-2 and Exp.3-4, it is evident that using the MemoryEQA framework leads to further improvements, indicating that the memory structure in MemoryEQA is more suitable for the EQA task. Furthermore, by comparing Exp.5-6 and Exp.9-12, it can be concluded that the Spatial LLM provides limited benefits for the EQA task. This may be due to the scarcity of questions related to spatial relationships (e.g., ”What is the size of the sofa?”) in the dataset, which prevents the Spatial LLM from fully exploiting its potential, resulting in its performance degrading to that of a standard LLM. It sum up, MemoryEQA exhibits superior performance in multi-modal reasoning tasks, particularly in complex scene understanding and knowledge integration.
The analysis of experimental results across Exp.3-4, Exp.7-8, and Exp.13-15 reveals a significant positive correlation between the performance of the VLM and the effectiveness of the EQA system. This observation underscores the critical role that VLM plays in enhancing the EQA system’s ability to process and interpret complex queries within visual environments. As the VLM’s accuracy and understanding improve, so does the EQA system’s capacity to deliver precise and contextually relevant answers, demonstrating a synergistic relationship between the two components. This finding highlights the importance of advancing VLM capabilities to further boost the overall performance of EQA systems in practical applications.
Table 3: Experiment on the ablation of memory strategies. (U: Memory update strategy. K: K-value selection strategy. R: Memory retrieval strategy.)
Strategy MT-HM3D HM-EQA
Succ. $\uparrow$ Step $\downarrow$ Succ. $\uparrow$ Step $\downarrow$
None 33.18 0.66 44.31 0.74
U 33.41 0.67 44.29 0.71
U + R 39.69 0.45 53.16 0.53
U + R + K 41.95 0.4 56.65 0.49
Table 4: Ablation study on MT-HM3D and HM-EQA to evaluate the impact of the memory information into different modules on performance. (S: Stop Criterion, A: Answering Module, P: Planner)
@@
None 30.22 0.64 51.55 0.75
S 35.1 0.64 52.24 0.7
S + A 40.99 0.61 54.76 0.71
S + A + P 41.95 0.4 56.65 0.49
Figure 3: Visualization of the exploration process. Memory plays a pivotal role in embodied question answering, enabling context-aware decision-making across various environments. $\texttt{RETREVAL}(\cdot)$ represents retrieving relevant memories from the memory library $\mathcal{M}$. $\texttt{UPDATE}(\cdot)$ represents updating the memory obtained from the current observation to the memory library $\mathcal{M}$. $\texttt{DECISION}(\cdot,\cdot)$ represents determining the next exploration direction based on memory and current observation. $\texttt{ANSWERING}(\cdot)$ represents answering based on memory. $m_{i}$ of the observation image represents the memory information under current observation $o_{i}$ (simplified version)